[Hallc_running] [New Logentry] Update on Hall-C Running

latif at jlab.org latif at jlab.org
Fri Jan 26 08:25:03 EST 2018

Logentry Text:
Finally Hall-C got back beam for a short time later in the owl shift. The target cool down started around 6 PM and ended around 1:00 PM. The Loop 3 fan stalled during first attempt of the condensation leading extended time to cool down (see https://logbooks.jlab.org/entry/3520973 for details), the cool down of LH2 was postponed to today. While attempting to deliver beam, MCC discovered gate valve was locked which required them to contact an off-site expert to come and open it (see https://logbooks.jlab.org/entry/3521053 for details). There is also on-going also down time for bleed-through from Hall-A tripping Hall-C ion chamber (Details: https://logbooks.jlab.org/entry/3521108). As of now we were waiting to get that resolved. The checklist suggested during yesterday's RC meeting was followed (see https://logbooks.jlab.org/entry/3520853) for the preparation of beam return.

We will have daily RC meeting at 1:30 PM today in the counting house conference room. Remote people can use the  bluejeans meeting ID: 543 903 672


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3521182
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180126/4c3b9bf9/attachment.html>

More information about the Hallc_running mailing list